OpenStudy (anonymous):

Find the solution of the inequality for 8

OpenStudy (anonymous):

not without starting with \[0<7x-x^2-8\]

OpenStudy (anonymous):

or better still \[x^2-7x+8<0\]

OpenStudy (dtan5457):

further adding, on simplify the quadratic and test the solutions,

OpenStudy (anonymous):

unfortunately this one does not factor you have to find the zeros using the the quadratic formula

OpenStudy (anonymous):

how do you do that??

OpenStudy (anonymous):

if you don't know the quadratic formula you cannot

OpenStudy (anonymous):

\[ax^2+bx+c=0\\ x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

oh I remember that now!

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

when you get the two zeros, since this is a parabola that opens up, it will be negative between the two and positive outside you want negative

OpenStudy (anonymous):

\[x^2-7x+8<0\] \[x^2-7x<-8\] \[x^2-7x+\left( \frac{ 7 }{ 2 } \right)^2<-8+\left( \frac{ 7 }{ 2 } \right)^2\] \[\left( x-\frac{ 7 }{ 2 } \right)^2<\frac{ -32+49 }{ 4 }\] \[\left| x-\frac{ 7 }{ 2 } \right|<\frac{ \sqrt{17} }{ 2 }\] \[-\frac{ \sqrt{17} }{ 2 }<x-\frac{ 7 }{ 2 }<\frac{ \sqrt{17} }{ 2 }\] adding 7/2 to each inequality \[\frac{ 7-\sqrt{17} }{ 2 }<x<\frac{ 7+\sqrt{17} }{ 2 }\]